OpenStudy (anonymous):

Partial Fractions Decomposition\[\frac{s-1}{[(s-1)^2+1](s^2-s)}\]

OpenStudy (anonymous):

\[\frac{As+B}{(s-1)^2+1}+\frac{C}{s}+\frac{D}{s-1}=\frac{s-1}{[(s-1)^2+1]s(s-1)}\]

OpenStudy (anonymous):

Is that a good start?

OpenStudy (anonymous):

\[(As+B)s(s-1)+C[(s-1)^2+1](s-1)+Ds[(s-1)^2+1]=s-1\]

OpenStudy (anonymous):

For s = 1:\[D=0.\]For s = 0:\[C=\frac{1}{2}.\]For s = -1:\[-2A+2B=3.\]For s = 2:\[2A+B=0.\]Therefore,\[B=1,\]\[A=-\frac{1}{2},\]and\[\frac{-1/2s+1}{(s-1)^2+1}+\frac{1/2}{s}=\frac{s-1}{[(s-1)^2+1]s(s-1)}.\]Is this correct?

OpenStudy (anonymous):

Rewriting:\[-\frac{1}{2}\frac{s}{(s-1)^2+1}+\frac{1}{(s-1)^2+1}+\frac{1}{2}\frac{1}{s}=\frac{s-1}{[(s-1)^2+1]s(s-1)}.\]

myininaya (myininaya):

thats exactly what i got but i went ahead in canceled the s-1's at the beginning

OpenStudy (anonymous):

it looks good

myininaya (myininaya):

\[\frac{1}{s(s^2-2s+2)}=\frac{A}{s}+\frac{Bs+D}{s^2-2s+2}=\frac{As^2-2As+2A+Bs^2+Ds}{s(s^2-2s+2)}\] =>\[1=s^2(A+B)+s(-2A+D)+(2A)\] => \[A+B=0; -2A+D=0; 2A=1 \]

myininaya (myininaya):

A=1/2 B=-1/2 D=1

OpenStudy (amistre64):

\[\frac{s-1}{((s-1)^2+1)(s^2-s)}\] \[\frac{s-1}{((s-1)^2+1)\ s(s-1)}\] \[\frac{1}{((s-1)^2+1)\ s}\] \[\frac{1}{s(s^2-2s+1)+s}\] \[\frac{1}{s^3-2s^2+s+s}\] \[\frac{1}{s^3-2s^2+2s}\] \[\frac{1}{s(s^2-2s+2)}\] \[s=\frac{2\pm\sqrt{-4}}{2}\]so its irreducible there across the reals \[\frac{1}{s(s^2-2s+2)}=\frac{C_1}{s}+\frac{C_2s+C_3}{s^2-2s+2}\]

myininaya (myininaya):

so awesome!

OpenStudy (amistre64):

how would you partial it across the complex and have it work out?

myininaya (myininaya):

why would you want to partial it across the complex?

OpenStudy (amistre64):

so that all you uners are linears

OpenStudy (amistre64):

denoms ... unders

OpenStudy (amistre64):

\[\frac{1}{s(s^2-2s+2)}=\frac{C_1}{s}+\frac{C_2}{s+(1+i)}+\frac{C_3}{s+(1-i)}\] or some such

OpenStudy (amistre64):

i spose those should be subtractions

OpenStudy (amistre64):

i tend to get lost in complexes, i just dont work with them enough